OpenStudy (istim):

Determine the projection of the first vector on the second. a=[6,-1], b=[11,5]

OpenStudy (istim):

I used the formula \[((a \times b)\div(b \times b)) \times b\]

OpenStudy (amistre64):

the vector projection? or its magnitude?

OpenStudy (istim):

MY final answer was (5.94,-1) Vector projection I think.

OpenStudy (istim):

Answer is [4.6,2.1]

OpenStudy (istim):

Correct answer.

OpenStudy (amistre64):

b.a ---- b |b|^2 since the naming is a bit backwards

OpenStudy (istim):

So I'm using the right formula?

OpenStudy (amistre64):

yes \[|a| cos(t)\frac{b}{|b|}\] cos * unit

OpenStudy (istim):

So how am I getting the wrong answer?

OpenStudy (istim):

Heres my work

OpenStudy (amistre64):

a=[6,-1] ; 36 1 = sqrt(37) b=[11,5] ; 121 25 = sqrt(146) ------ 66-5 = 61 61 ---- <11,5> 146

OpenStudy (istim):

=(((6,-1)*(11,5))/((11,5)(11,5)))(11,5) =((6*11,-1,5)/(11*11,5*5))(11,5) =((66,-5)/(121,25))(11,5) =(0.54,0.2)(11,5) =(5.94,-1)

OpenStudy (istim):

Wait. What are you doing in your steps? IT's a little confusing.

OpenStudy (amistre64):

finding my parts the under is a.b and to each side is the magnitudes of each

OpenStudy (amistre64):

magnitude is: square add and sqrt

OpenStudy (istim):

Your lingo is confusing.

OpenStudy (amistre64):

=((66,-5)/(121,25))(11,5) =(0.54,0.2)(11,5) <---- .54,0.2 what is that? =(5.94,-1)

OpenStudy (amistre64):

both parts of the vector is positive; and your scalar is positive; how do you end up with a - anything in the end result?

OpenStudy (istim):

0.54 is 66/121 0.2 is -5/25

OpenStudy (amistre64):

ok, but what is that? \[\frac{66-5}{121+25}=\frac{61}{146}\]

OpenStudy (amistre64):

youre fine up till you decide to invent your own thing there

OpenStudy (istim):

I dunno. What is that? I didn't write that. I wrote ((66,-5)/(121,25))

OpenStudy (amistre64):

i know what you wrote; and you wrote some magical fantasy spell that has nothing to do with finding what you started out to find in the first place

OpenStudy (amistre64):

66 and -5 are components of a dot product; you combine them to get a single value .... as defined by a dot product. the magnitude of b is a single value, not some ordered pair ...

OpenStudy (istim):

But the final answer is [4.6,2.1]...

OpenStudy (istim):

According to da txtbook.

OpenStudy (amistre64):

yes it is; which is: \[\frac{61}{146}<11,5>\implies <\frac{61(11)}{146},\frac{61(5)}{146}>\]

OpenStudy (amistre64):

this decimal out to about 4.6 and 2.1

OpenStudy (istim):

Ok. Lemme try it out. I guess you shoulddo more importantthings as I write it down.

OpenStudy (amistre64):

well, i DO have to urinate :)

OpenStudy (istim):

I'm done. Thank you very much, amistre64.

OpenStudy (amistre64):

youre welcome :)